# VisFly: An Efficient and Versatile Simulator for Training Vision-based Flight

### v1

@@ Abstract
State-of-art simulators primarily focus on providing full-stack simulation tools or state-only parallelizability. Due to the limitation of computing resources, they have to make trade-off among photo-realism and sampling efficiency. Yet, both factors are crucial for data-driven reinforcement learning tasks. Therefore, we introduce a both rapid-rendering and photo-realistic quadrotor simulator: VisFly. VisFly offers a user-friendly framework and interfaces for users to develop or utilize. It couples differentiable dynamics and habitat-sim rendering engines, reaching frame rate of up to 10000 frame per second in cluttered environments. The simulation is wrapped as a gym environment, facilitating convenient implementation of various baseline learning algorithms. It can directly import all the open-source scene datasets compatible with habitat-sim, which provides more fair benchmarks for comparing the intelligent policy. VisFly presents a general policy architecture for tasks, and the whole framework is verified by three regular quadrotor tasks with visual observation. We will make this tool available at https://github.com/SJTU-ViSYS/VisFly.
@@
Quadrotor Simulation, Vision-based Learning, Gym-wrapped Framework, Benchmark Visual Tasks.
@@ I Introduction
Recently, researchers have demonstrated that reinforcement learning (RL) facilitates not only high-level decision-making applications such as autonomous driving [8], strategy games[9], but also basic control problems for complicated systems, like humanoid [10], quadruped locomotion [locomotion]. These potential and powerful applications inspire academics in the field of aerial robotics to explore reinforcement learning to enhance the autonomous flight [11, 12]. RL-based approach usually relies a simulator to collecting massive data generated from the interaction between the agent and the simulated environment. There are already well-designed simulators [13, 14] widely used in the robotics community, with some specifically dedicated to quadrotors [15, 16]. These simulators are primarily designed for fast prototyping and iteration, typically encompassing the full stack of hardware and software simulations with an emphasis on closed-loop validation and testing. They usually operate at slow frame rates (10 to 100 fps), which does not negatively impact testing or validation but presents significant drawbacks for reinforcement learning due to low sampling efficiency. Some quadrotor simulators [17, 18, 19] have addressed these drawbacks by improving the sampling of physical state data, enabling the training of state-based policies for autonomous flight. However, they still lack the capability to sample vision inputs at high frame rates, making it challenging to train a high-performance vision-based flight policy.
Figure 1: Overall diagram of VisFly. The differentiable physics and Habitat-sim rendering engine constitute the quadrotor agents, which is then wrapped as gym environments, preserving interfaces for RL algorithms.
@@
Given that vision is one of the most informative and cost-effective sources for perception, it holds substantial potential for deploying a vision-based autonomous flight model in real-world applications. The key to effectively training a vision-based flight policy for quadrotors via reinforcement learning lies in 1) the rendering speed of the vision data, 2) efficient computation of flight physics, 3) the diversity of scenes for training, and 4) the adoption of state-of-the-art reinforcement learning algorithms. Moreover, to enable the autonomous flight of a team of quadrotors, multi-agent learning should also be supported.
@@
To achieve these goals, we developed an efficient simulation environment for training vision-based autonomous flight of quadrotors. Our environment leverages the PyTorch, Habitat-Sim [20], and OpenAI’s Gym [21] platforms. We implemented flight dynamics and three widely-used flight controllers using PyTorch, making flight physics not only computationally efficient but also differentiable, thereby enabling the use of model-based algorithms with known dynamics [22]. By adopting Habitat-Sim’s rendering engine, our environment can achieve high frame rates at 10,000 frames per second while rendering color images at a resolution of 64×64, which is about two orders of magnitude faster than existing quadrotor simulators. Moreover, our simulator can load a series of open-source datasets comprising a vast number of 3D scene models (including real-world 3D scans), allowing us to train vision-based flight policies in diverse and realistic scenes. Our simulator is highly flexible, featuring Gym-wrapped interfaces for integrating the latest learning algorithms and customized tasks. With our well-designed parallel mechanism, we can run multiple agents within different scenes simultaneously, facilitating training vision-based flight policies in multi-agent settings.
@@
To validate our environment, we present an example of using a general network architecture to train vision-based flight policies for three autonomous flight tasks: landing, navigation, and cooperative flight through narrow gaps. All tasks rely solely on vision sensory inputs without any external guidance. The results demonstrate that our simulation environment can be used to train vision-based flight policies in both single and multi-agent settings with high success rates in a very short period. Our simulation environment will be released as an open-source project and will be continuously updated. To summarize, the contributions of this work include:
@@
    We present a novel quadrotor simulator that achieves high performance of rendering both physics and vision data, allowing to train vision-based flight policies efficiently.
@@
Table I: Key Character Comparison of Popular Simulators
Engine Parallel FPS Sensor Gym- 3D Differ- Swarm
Physics Render agent scene Physics Render wrap Datasets entiable
RotorS Gazebo OpenGL $1\times 10^{2}$ 20 I, D, C
Airsim PhysX Unreal $1\times 10^{2}$ 60 I, D, C, S
Flightmare Flexible Unity $2.2\times 10^{5}$ $2.3\times 10^{2}$ I, D, C, S
PyBulletDrone Bullet OpenGL $1\times 10^{4}$ $3\times 10^{3}$ I, D, C, S
FlightGoggles Flexible Unity $1\times 10^{3}$ 64 I, D, C, S
CrazyS Gazebo OpenGL $1\times 10^{3}$ 20 I, D, C
OmniDrone Self-defined Isaac Sim $2\times 10^{5}$ I, D, C, S
FastSim Flexible Unity $1\times 10^{2}$ 40 I, D, C, S
Ours Self-defined Habitat-sim $3.7\times 10^{4}$ $1\times 10^{4}$ I, D, C, S
  • 1
    I, C, D, S refer to IMU, RGB, depth, segmentation. The simulators with superscript * are full-stack, while the others are learning-specialized. VisFly is light-weight, which does not have too much requirements for hardware. Noting that, FPS data provided by original researches are not evaluated with same number of agents, but it still reflect the performance to a large extent. ”-” in FPS column means this function is unavailable. All the frame rates are test with 100 environments if parallelization is available. Render FPS are test while rendering $64\times 64$ depth images. If the simulator cannot be deployed locally at present, FPS will be directly cited from original paper.
@@ II-A Full-stack simulators
The initial aim of simulators is to create a simulated platform for users to verify the algorithmic prototypes because quadrotors with immature algorithm are easy to crash. This category always consists of comprehensive full-stack simulated flight modules such as perception, planning, and control, especially for highly realistic physics. Gazebo [13] is a widely used full-stack robotics simulator of this category. While it is general-purpose simulator, it has strong support for UAVs through various plugins and models. Considering various requirements, researchers develop simulators specific for UAVs based on Gazebo. RotorS [15] is a popular multi-rotor UAV modular simulator built on Gazebo, provides flexible interfaces to replace several parts of the pipeline, making it convenient to verify state estimators or control algorithms. Rather than using open-source rendering engines as in Gazebo, several simulators utilize commercial rendering engines such as Unity  [^1] https://unity.com/ and Unreal Engine  [^2] https://www.unrealengine.com/ to achieve high visual fidelity. For instance, FlightGoogles[23] employs the Unity engine for realistically simulating image sensors and uses real-world flight data to create avatars in simulation environments, enabling vehicle-in-the-loop simulation. Similarly, AirSim [14] is built on Unreal Engine to offer physically and visually realistic environments, along with real-time hardware-in-the-loop (HITL) simulation capabilities. Additionally, AvoidBench [24], built on the Unity engine, offers valuable benchmarks to evaluate obstacle avoidance algorithms. FastSim [25], also built on Unity, stands out as the first to integrate path planning algorithms directly into the simulator, facilitating tests of high-level autonomous control functionalities. While using commercial engines significantly enhances image quality (including depth maps and semantic segmentation), it comes at the cost of slower rendering speed because of their sophisticated rendering pipelines, particularly for the vision data. However, for applications focusing on learning in simulators using vision sensors, it is critical to acquire vision data quickly for fast training.
@@ II-B Learning-specialized simulators
High acquisition rates for both sensory data and robot physics significantly accelerate the training process of learning-based algorithms. This makes them essential features for simulators designed for learning. Notably, physics computation can be achieved at very high frequencies in some existing simulators. For instance, by paralleled running, the frame rate of physics computation can reach an impressive 100000 FPS in Flightmare [17] and OminiDrone [18]. However, the frame rates will drop down to hundreds of FPS if the visual observation is rendered simultaneously in Flightmare. While PyBulletDrone [19] achieves high rendering FPS, it focuses more on scene visualization rather than sensor data acquisition. Additionally, the basic OpenGL-based rendering engine (TinyRenderer for cpu rendering or OpenGL3 for GPU rendering ) lacks the functionality to import complex scenes and objects, making it difficult for training navigation policies using diverse and complicated scene data.
@@
Our simulator is built on Habitat-sim, utilizing its fast rendering engine and well-designed interfaces to leverage existing massive 3D datasets. However, unlike Habitat-sim, which focuses on ground robots and manipulations, our simulator is tailored specifically for aerial robots. For the convenience of implementation, existing learning-specialized simulators, such as Flightmare, OminiDrone, PyBulletDrone wrap the environment upon OpenAI’s gym environment standards. An OpenAI’s gym class receives an action and updates its internal state, and returns a new state paired with a corresponding reward. [TODO] but the examples are highly customized for specific tasks. Each of them provides individual environments or scenes, making it hard to evaluate the performance at a common and fair metric. Therefore, we also adopt a modular design for integrating the Gym framework in our simulator to make it easier to customize the vision-based flight tasks. To highlight the novel features of our simulator, we compare it with existing simulators in Table I.
@@ III Methodology
VisFly consists of three parts: (1) Differentiable, and parallelized dynamics. (2) Rapid-rendering, parallelized, photo-realistic scene manager built by habitat-sim. (3) Gym-wrapped interfaces and its extensions. A schematic diagram of VisFly is shown in Figure 1.
Figure 2: RGB, depth, segmentation photos captured in HM3D, MP3D, HSSD, Replica, and customized scenes by generation. The right subfigure is the swarm of quadrotors in a clear garage.
III-A Dynamics
The dynamics simulation is implemented using PyTorch, allowing physics to run on GPUs. Furthermore, PyTorch’s computational graph enables the computation of analytical gradients through backpropagation, which makes it a differentiable process. It could be directly leveraged as a known dynamic transition model in model-based RL [26]. Analytical gradients can descend along much more precise directions because it follows the precise dynamic process, thereby accelerating training several times compared to common value-based or policy-based reinforcement learning [22, 27]. Considering general multi-rotors dynamics, we build the quadrotor model as:
@@ III-A Dynamics
where $\mathbf{x}_{W}$, $\mathbf{v}_{W}$, $\mathbf{q}$ denote position, velocity, orientation (quaternion) of quadrotors in world coordinate, and $\mathbf{\Omega}$ denotes angular velocity in body coordinate. $\otimes$ is multiplication between quaternion and vector. $\mathbf{R}_{WB}$ is the matrix of transformation from body coordinate to world coordinate. $\mathbf{f}$, $\mathbf{\eta}$ denote the collective rotor thrust and momentum. $\mathbf{J}$ indicates the diagonal inertia matrix, while $m$ indicates the quadrotor mass. $\mathbf{d}$ represents the air drag in the body coordinate, assumed positive with the square of velocity [28]:
$\mathbf{d}=\frac{1}{2}\rho\mathbf{v}_{B}\odot\mathbf{v}_{B}\mathbf{C}_{d}\odot\mathbf{s}$ (2)
@@
where $\rho$, $\mathbf{C}_{d}$, $\mathbf{s}$, $\mathbf{v}_{B}$ denote air density, drag coefficients, crossing area, velocity in body coordinates. $\odot$ refers to element-wise multiplication. $\mathbf{f}$, $\mathbf{\eta}$ are computed from single rotor thrust $\mathbf{f}_{i}$:
$\mathbf{f}=\sum_{1}^{4}f_{i}\qquad\mathbf{\eta}=\sum_{1}^{4}\mathbf{T}_{i}\times\mathbf{f}_{i}$ (3)
@@
$\mathbf{T}_{i}$ denotes the moment arm matrix, with $i$ representing the rotor index. To account for control input delay, we introduce the time decay constant $c$ and model it as an exponential process:
$\mathbf{f}_{i}(3)=k_{2}\omega_{i}^{2}+k_{1}\omega_{i}+k_{0}\qquad\omega_{i}=\omega_{i}^{des}+(\omega_{i}^{\prime}-\omega_{i}^{des})e^{-ct}$ (4)
@@
where $\omega$ denotes the rotor speed, and superscripts , $des$ separately indicate current and desired state. $k_{2,1,0}$ are constant thrust coefficients w.r.t. rotor speed. Based on [29, 30], we develop three widely-used control methods, individual single-rotor thrusts (SRT), mass-normalized collective thrust and bodyrates (CTBR), position commands and yaw (PS), and linear velocity commands and yaw (LV). SRT and CTBR are end control inputs for Betaflight, which can be used for training efficient end2end policies[31]. The minimal simulation time step can be set several times shorter than the control time step, resulting in more precise physics. Both 4th order Runge-Kutta and Euler methods are supported, providing users the flexibility to choose based on their platform’s performance.
III-B Scenes

### v3

@@ Abstract
We present VisFly, a quadrotor simulator designed to efficiently train vision-based flight policies using reinforcement learning algorithms. VisFly offers a user-friendly framework and interfaces, leveraging Habitat-Sim’s rendering engines to achieve frame rates exceeding 10,000 frames per second for rendering motion and sensor data. The simulator incorporates differentiable physics and seamlessly integrates with the Gym environment, facilitating the straightforward implementation of various learning algorithms. It supports the direct import of all open-source scene datasets compatible with Habitat-Sim, enabling training on diverse real-world environments and ensuring fair comparisons of learned flight policies. We also propose a general policy architecture for three typical flight tasks relying on visual observations, which have been validated in our simulator using reinforcement learning. The simulator will be available at [https://github.com/SJTU-ViSYS/VisFly].
@@
Quadrotor Simulation, Vision-based flight, learning-based navigation.
@@ I Introduction
Table I: Key Character Comparison of Popular Simulators
Engine Parallel FPS Sensor Gym- 3D Different- Swarm
Physics Render agent scene Physics Render wrap Datasets iable
RotorS[4] Gazebo OpenGL $1\times 10^{2}$ 20 I, D, C
Airsim[5] PhysX Unreal $1\times 10^{2}$ 60 I, D, C, S
FlightGoggles[6] Flexible Unity $1\times 10^{3}$ 64 I, D, C, S
FastSim [7] Flexible Unity $1\times 10^{2}$ 40 I, D, C, S
CrazyS [8] Gazebo OpenGL $1\times 10^{3}$ 20 I, D, C
Flightmare[9] Flexible Unity $2.2\times 10^{5}$ $2.3\times 10^{2}$ I, D, C, S
PyBulletDrone[10] Bullet OpenGL $1\times 10^{4}$ $3\times 10^{3}$ I, D, C, S
OmniDrone[11] Self-defined Isaac Sim $2\times 10^{5}$ $3\times 10^{2}$ I, D, C, S, L
Ours Self-defined Habitat-sim $3.7\times 10^{4}$ $1\times 10^{4}$ I, D, C, S
  • 1
    I, C, D, S, and L refer to IMU, RGB, depth, segmentation, and LiDAR, respectively. Simulators marked with an asterisk (*) are full-stack, while the others are learning-specialized. VisFly is lightweight and does not have high hardware requirements. Note that the FPS data provided by the original papers were not evaluated with the same number of agents, but they still offer a performance reference. A dash (”-”) in the FPS column indicates that this function is unavailable. All frame rates are tested with 100 environments if parallelization is supported, and render FPS are tested while rendering $64\times 64$ depth images. FastSim cannot be deployed locally at present, so the FPS data is directly cited from the original paper. All frame rates are tested on a desktop with an RTX 4090 and a 32-core 13th Gen Intel(R) Core(TM) i9-13900K.
@@
Recently, researchers have demonstrated that reinforcement learning (RL) facilitates not only high-level decision-making applications such as autonomous driving [16], strategy games[17], but also basic control problems for complicated systems, like humanoid [18], quadruped locomotion [19]. These potential and powerful applications inspire academics in the field of aerial robotics to explore reinforcement learning to enhance autonomous flight [20, 21]. The RL-based approach usually relies on a simulator to collect massive data generated from the interaction between the agent and the simulated environment. There are already well-designed simulators [22, 5] widely used in the robotics community, with some specifically dedicated to quadrotors [4, 8]. These simulators are primarily designed for fast prototyping and iteration, typically encompassing the full stack of hardware and software simulations with an emphasis on closed-loop validation and testing. They usually operate at slow frame rates (100 fps), which does not negatively impact testing or validation but presents significant drawbacks for reinforcement learning due to low sampling efficiency. Some quadrotor simulators [9, 11, 10] have addressed these drawbacks by improving the sampling of physical state data, enabling the training of state-based policies for autonomous flight. However, they still lack the capability to sample vision input at high frame rates, making it challenging to train a high-performance vision-based flight policy.
@@
Given that vision is one of the most informative and cost-effective sources for perception, it holds substantial potential for deploying a vision-based autonomous flight model in real-world applications. The key to effectively training a vision-based flight policy for quadrotors via reinforcement learning lies in 1) the rendering speed of the vision data, 2) the efficient computation of flight physics, 3) the diversity of scenes for training, and 4) the adoption of state-of-the-art reinforcement learning algorithms. Moreover, to enable the autonomous flight of a team of quadrotors, multi-agent learning should also be supported.
@@
To achieve these goals, we developed an efficient simulation environment for training vision-based autonomous flight of quadrotors. Our environment leverages the PyTorch, Habitat-Sim [23], and OpenAI’s Gym [24] platforms. We implemented flight dynamics and three widely-used flight controllers using PyTorch, making flight physics not only computationally efficient but also differentiable, thereby enabling the use of model-based algorithms with known dynamics [25]. By adopting Habitat-Sim’s rendering engine, our environment can achieve high frame rates at 10,000 frames per second while rendering color images at a resolution of 64×64, which is about two orders of magnitude faster than existing traditional quadrotor simulators. Moreover, our simulator can load a series of open-source datasets comprising a vast number of 3D scene models (including real-world 3D scans), allowing us to train vision-based flight policies in diverse and realistic scenes. Our simulator is highly flexible, featuring gym-wrapped interfaces for integrating the latest learning algorithms and customized tasks. With our well-designed parallel mechanism, we can run multiple agents within different scenes simultaneously, facilitating training vision-based flight policies in multi-agent settings.
@@
To validate our environment, we present an example of using a general network architecture to train vision-based flight policies for three autonomous flight tasks: landing, navigation, and cooperative flight through narrow gaps. All tasks rely solely on self-provided information, including vision and sensory inputs, without any external guidance. The results demonstrate that our simulation environment can be used to train vision-based flight policies in both single and multi-agent settings with high success rates in a very short period. Our simulation environment will be released as an open-source project and will be continuously updated. To highlight the novel features of our simulator, we compare it with existing simulators in Table I. To summarize, the contributions of this work include:
@@
    We present a novel quadrotor simulator that achieves high performance in rendering both physics and vision data, allowing to train vision-based flight policies efficiently.
@@ II-A Full-stack simulators
We compare the key factors of primary simulators in Table I, which provides a quick overview of the progress. The initial aim of simulators is to create a simulated platform for users to verify the algorithmic prototypes because quadrotors with immature algorithms are easy to crash. This category always consists of comprehensive full-stack simulated flight modules such as perception, planning, and control, especially for highly realistic physics. Gazebo [22] is a widely used full-stack robotics simulator of this category. While it is a general-purpose simulator, it has strong support for UAVs through various plugins and models. Considering various requirements, researchers develop simulators specific for UAVs based on Gazebo. RotorS [4] is a popular multi-rotor UAV modular simulator built on Gazebo, providing flexible interfaces to replace several parts of the pipeline, making it convenient to verify state estimators or control algorithms. CrazyS [8] is a lightweight simulator specialized for a software-to-hardware solution Crazyflie 2.0, which helps to easily deploy in the real world. Rather than using open-source rendering engines as in Gazebo, several simulators utilize commercial rendering engines such as Unity  [^1] https://unity.com/ and Unreal Engine  [^2] https://www.unrealengine.com/ to achieve high visual fidelity. For instance, FlightGoogles[6] employs the Unity engine for realistically simulating image sensors and uses real-world flight data to create avatars in simulation environments, enabling vehicle-in-the-loop simulation. Similarly, AirSim [5] is built on Unreal Engine to offer physically and visually realistic environments, along with real-time hardware-in-the-loop (HITL) simulation capabilities. Additionally, AvoidBench [26], built on the Unity engine, offers valuable benchmarks to evaluate obstacle avoidance algorithms. FastSim [7], also built on Unity, stands out as the first to integrate path planning algorithms directly into the simulator, facilitating tests of high-level autonomous control functionalities. While using commercial engines significantly enhances image quality (including depth maps and semantic segmentation), it comes at the cost of slower rendering speed because of their sophisticated rendering pipelines, particularly for the vision data. However, for applications focusing on learning in simulators using vision sensors, it is critical to acquire vision data quickly for fast training.
@@ II-B Learning-specialized simulators
High acquisition rates for both sensory data and robot physics significantly accelerate the training process of learning-based algorithms. This makes them essential features for simulators designed for learning. Notably, physics computation can be achieved at very high frequencies in some existing simulators. For instance, by paralleled running, the frame rate of physics computation can reach an impressive 100000 FPS in Flightmare [9] and OmniDrone [11]. However, the frame rates will drop down to hundreds of FPS if the visual observation is rendered simultaneously in Flightmare. PyBulletDrone [10] achieves high rendering FPS via simple rendering engine OpenGL3 and TinyRender  [^3] https://github.com/ssloy/tinyrenderer. However, the direct basic OpenGL-based rendering engine lacks the functionality to import complex scenes and objects, making it difficult to train navigation policies using diverse and complicated scene data.
@@
Our simulator is built on Habitat-sim, utilizing its fast rendering engine and well-designed interfaces to leverage existing massive 3D datasets. However, unlike Habitat-sim, which focuses on ground robots and manipulations, our simulator is tailored specifically for aerial robots. For the convenience of implementation, existing learning-specialized simulators, such as Flightmare, OmniDrone, and PyBulletDrone wrap the environment upon OpenAI’s gym environment standards. An OpenAI’s gym class receives an action, updates its internal state, and returns a new state paired with a corresponding reward. Like most of existing simulators, VisFly provides examples and tutorials to help beginners get started quickly. However, unlike others that primarily focus on state-only tasks—using only the UAV’s state, such as orientation, position, and velocities for control or planning—VisFly’s examples are specifically designed for vision-based flights. While some simulators, like OmniDrone, include examples using LiDAR inputs, VisFly emphasizes using visual data. Our examples are released with a modular design and are seamlessly integrated with Gym environments, providing users with an easy way to customize their vision-based flight tasks.
Figure 1: Overall diagram of VisFly. Our system uses differentiable physics to drive single or multiple quadrotor agents through four different types of controllers and employs Habitat-sim’s rendering engine for fast rendering and access to open-source datasets. All of these components are integrated into Gym environments, providing standard interfaces for various learning algorithms.
@@ III Methodology
Here, we introduce three key features of VisFly and how they are implemented: (1) differentiable, and parallelized dynamics computations; (2) a rapid-rendering, parallelized, photo-realistic scene manager; (3) Gym-wrapped interfaces and extensions based on modular designs. A schematic diagram of VisFly is shown in Figure 1.
III-A Differentiable dynamics based on PyTorch
The dynamics simulation is implemented using PyTorch to enable efficient, parallel physics computations on GPUs. PyTorch’s computational graph also enables the computation of analytical gradients through back-propagation, making the dynamics computation differentiable. The differentiable dynamic model can be directly integrated into model-based reinforcement learning frameworks[27]. By following the exact dynamic process, analytical gradients offer more precise descent directions, significantly accelerating training compared to common value-based or policy-based reinforcement learning methods[25, 28]. We model quadrotor’s dynamics as follows:
@@ III-A Differentiable dynamics based on PyTorch
where $\mathbf{x}_{W}$, $\mathbf{v}_{W}$, $\mathbf{q}$ represent the position, velocity, and orientation (quaternion) of quadrotors in the world frame, and $\mathbf{\Omega}$ denotes the angular velocity in the body frame. The quaternion-vector product is represented by $\otimes$. The rotation matrix from the body frame to the world frame is $\mathbf{R}_{WB}$. The collective rotor thrust and the momentum are denoted by $\mathbf{f}$ and $\mathbf{\eta}$, respectively. The diagonal inertia matrix is $\mathbf{J}$, and the quadrotor mass is $m$. The air drag in the body coordinate, assumed proportional to the square of velocity [29], is represented by $\mathbf{d}$ :
$\small\mathbf{d}=\frac{1}{2}\rho\mathbf{v}_{B}\odot\mathbf{v}_{B}\mathbf{C}_{d}\odot\mathbf{s}$ (2)
@@
where $\rho$, $\mathbf{C}_{d}$, $\mathbf{s}$, $\mathbf{v}_{B}$ denote the air density, drag coefficients, crossing area, and velocity in the body frame, respectively, the element-wise multiplication is denoted by $\odot$. The collective thrust and momentum, $\mathbf{f}$ and $\mathbf{\eta}$, are computed from individual rotor thrusts, $\mathbf{f}_{i}$:
$\small\mathbf{f}=\sum_{1}^{4}f_{i},\quad\mathbf{\eta}=\sum_{1}^{4}\mathbf{T}_{i}\times\mathbf{f}_{i}.$ (3)
@@
Here $\mathbf{T}_{i}$ denotes the moment arm matrix for the rotor $i$. To account for control input delay, we introduce the time decay constant $c$ and model it as an exponential process:
$\small\mathbf{f}_{i}(3)=k_{2}\omega_{i}^{2}+k_{1}\omega_{i}+k_{0},\quad\omega_{i}=\omega_{i}^{des}+(\omega_{i}^{\prime}-\omega_{i}^{des})e^{-ct}$ (4)
@@
where $\omega_{i}$ denotes the spinning speed for rotor $i$, and $\omega_{i}^{\prime}$, $\omega_{i}^{des}$ indicate the current and desired states, respectively, the constants $k_{2},k_{1},k_{0}$ are thrust coefficients with respect to rotor speeds. Based on [30, 31], we develop three widely-used control methods: individual single-rotor thrusts (SRT), mass-normalized collective thrust and body rates (CTBR), position commands and yaw (PS), and linear velocity commands and yaw (LV). SRT and CTBR are control inputs for Betaflight, which can be used for training efficient end-to-end policies[32]. The minimal simulation time step can be set several times shorter than the control time step, resulting in more precise physics. Both the fourth-order Runge-Kutta and Euler methods are supported, providing users the flexibility to choose based on their platform’s performance.
Figure 2: RGB, depth, and semantic segmentation captured in open-source and customized scenes. On the right is a swarm of quadrotors in a clear garage.
III-B Efficient rendering and open-source scene management